An exercise machine with an original price of $652 is on sale at 20% off. What is the exercise machine's sale​ price?
The exercise​ machine's sale price is ​$??. ​(Round to the nearest cent as​ needed.)
Suppose 22 pints of aa 21​% alcohol solution is mixed with 8 pints of an 80​% alcohol solution. What is the concentration of alcohol in the new 10​-pint ​mixture?
The concentration of alcohol in the new 10​-pint mixture is ?

Answer by stanbon(75887)

An exercise machine with an original price of $652 is on sale at 20% off. What is the exercise machine's sale​ price?
Equation::
Sale price = 652-0.2*652 = 0.8*652 = $521.60
-----------------------------------------------
The exercise​ machine's sale price is ​$??. ​(Round to the nearest cent as​ needed.)
---------------------------------
Suppose 2 pints of a 21​% alcohol solution is mixed with 8 pints of an 80​% alcohol solution. What is the concentration of alcohol in the new 10​-pint ​mixture?
Equation:
alch + alch = alch
0.21*2 + 0.8*8 = x*10
-------
0.42 + 6.4 = 10x
----
10x = 6.82
x = 0.682
---
Ans:: 68.2% (Percent alcohol in the final solution)
------------------------
The concentration of alcohol in the new 10​-pint mixture is ?
-----------
Cheers,
Stan H.
-------------